Originally Posted by James Lacey
Your 1 in 10,000 claim is wildly wrong. All your other blabber is just a distraction.
By the way, anyone who cares, and there aren't many, can determine how to convert percentages to odds and decide for themselves. I'm comfortable with my answer, and I know that 1 in 10,000 is wildly incorrect.
WARNING - THIS IS TOO DAMN LONG, DO NOT READ. (I worked this out for my own amusement and a math refresher course).
Rich is confusing the concept of probability with the concept of odds. Rich will
never admit this, but you are quite correct in your assessment of the odds, James. A good discussion of these concepts can be found here: https://journalfeed.org/article-a-day/2018/idiots-guide-to-odds-ratios/
For my own amusement, I worked it out. I know that you know this stuff. I had to look it up. In context it was clear that you listed a
probability of getting killed or severely injured in a knife fight as 99.99% (I know you were just exaggerating to make a point, we went through that yesterday). Probably being too pedantic, but for Rich's sake 99.99% = .9999. This is a
probability. This means that the probability of emerging unscathed from the knife fight is 1.0000 - 0.9999 = 0.0001.
Rich interjected the concept of "odds" and I know that you are just busting his balls "Rich style" by seizing on his mis-use of the word "odds" when he meant "probability" But hey, what would Rich do in similar circumstances? Like I said above, you played him like a maestro.
Here is where it gets interesting from a math perspective. "Odds" is a ratio, defined as the probability of something happening over the probability of that something not happening. https://sphweb.bumc.bu.edu/otlt/mph-modules/bs/bs704_confidence_intervals/BS704_Confidence_Intervals10.html#:~:text=The%20odds%20are%20defined%20as,not%20occurring%20is%201%2DY.
The example using playing cards in the first link above is quite informative. The probability of drawing a spade from a deck of cards is 13/52 = .25 or 25%. Probability is very intuitive. The odds of drawing a spade are much less intuitive.
The
odds in the card example are the ratio of the probability of drawing a spade over the probability of not drawing a spade.
Specifically, the
odds of drawing a spade are: .25/.75 = 1/3 pronounced 1
to 3 odds. (meaning you will overall draw 1 spade
to 3 other suits) Here is where I think most people get confused. They think the odds are pronounced 1
in 3, and that the odds of drawing a spade should intuitively be 1
in 4, not 1 in 3. What reasonable people have done there is ignored the mathematical definition of "odds" and converted back to the "probability" of drawing a spade in their heads since 1/4 is in fact the correct probability.
So, the
odds of
getting killed or maimed in your admittedly exaggerated knife attack are .9999/.0001 = 1 to 9999 - you were exactly correct, except for the original exaggeration and adding a decimal place later.
Jim